Problem 5-3 Future Value and Multiple Cash Flows [LO 1] Booker, Inc., has identified an investment project with the following cash flows. Year Cash Flow 1 $ 960 2 1,190 3 1,410 4 2,150 If the discount rate is 9 percent, what is the future value of these cash flows in Year 4? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Future value $ What is the future value at an interest rate of 12 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Future value $ What is the future value at an interest rate of 23 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Future value $
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.A=960*(1.09)^3+1190*(1.09)^2+1410*(1.09)^1+2150

=$6343.97(Approx).

b.A=960*(1.12)^3+1190*(1.12)^2+1410*(1.12)^1+2150

=$6570.67(Approx).

c.A=960*(1.23)^3+1190*(1.23)^2+1410*(1.23)^1+2150

=$7471.08(Approx).